What volume of 0.1 M NaOH is needed to titrate 100 mL of 0.2 M H2SO4?

a) 50 mL

b) 100 mL

c) 200 mL

d) 400 mL
1
Expert's answer
2014-06-17T06:32:20-0400
2NaOH + H2SO4 --> Na2SO4 + 2H2O
n(NaOH) = 2n(H2SO4), => c(NaOH)V(NaOH) = 2c(H2SO4)V(H2SO4)
V(NaOH) = 2c(H2SO4)V(H2SO4)/c(NaOH) = (2 * 0.2 M * 100 mL) / 0.1 M = 400 mL
So, the correct option is (d): 400 mL